# Node-like as a Whole: Structure-aware Searching and Coarsening for Graph Classification

### v2

@@ 5.1. Experimental Settings
We evaluate our approach on eight widely used datasets from TUDataset (Morris et al., 2020), including three social network datasets: COLLAB (CO), IMDB-BINARY (IB), and IMDB-MULTI (IM), two molecules datasets: NCI1 (N1) and NCI109 (N109), and three bioinformatics datasets: D&D (DD), PTC_MR (PTC), and PROTEINS (PRO). Table 1 shows the details of the datasets.
  • COLLAB is a scientific collaboration dataset in physics. A graph corresponds to a researcher’s ego network, which has three possible labels indicating the fields that the research belongs to.
  • IMDB-BINARY (MULTI) shows a movie collaboration network between actors/actresses and movies in IMDB. An edge connects two people when they participate in the same scene. The labels of BINARY are Action and Romance, and MULTI’s are Comedy, Romance, and Sci-Fi.
  • NCI1 (109) is relative to anti-cancer screens. Each input graph represents a chemical compound, treating atoms and bonds as nodes and edges. Classification tasks reveal whether the chemicals appear positive or not. NCI1 aims for non-small cells and NCI109 for ovarian cells.
  • D&D is a dataset of proteins based on the same data as PROTEINS, where nodes represent secondary structure elements with their type information and edges show spatial proximity.
  • PTC_MR reports the carcinogenicity for rats of 344 chemical compounds.
  • PROTEINS is a dataset of proteins. In each graph, nodes represent the amino acids, and edges exist if two nodes are less than 6 Angstroms apart. It includes two categories: enzymes and non-enzymes.
Table 2. The main results of the graph classification task on eight datasets (mean accuracy (%) and standard deviation). The best scores are in bold and underline. ’-’ indicates that results are unavailable in the original or published papers. A.R shows the average ranking of each method. We highlight every method with outstanding performance in each category.
Method CO IB IM DD N1 PTC PRO N109 A.R
GW $72.84^{0.28}$ $65.87^{0.98}$ $43.89^{0.38}$ $78.45^{0.26}$ $62.30^{0.30}$ $57.26^{1.41}$ $71.67^{0.55}$ 18.71
WL $79.02^{1.77}$ $73.40^{4.63}$ $49.33^{4.75}$ $79.78^{0.36}$ $\underline{\bf{82.19}}^{0.18}$ $57.97^{0.49}$ $74.68^{0.49}$ $\underline{\bf{82.46}}^{0.24}$ 8.75
GCN $81.72^{1.64}$ $73.30^{5.29}$ $51.20^{5.13}$ $79.12^{3.07}$ $76.00^{0.90}$ $59.40^{10.30}$ $75.65^{3.24}$ $67.09^{3.43}$ 9.13
GAT $75.80^{1.60}$ $70.50^{2.30}$ $47.80^{3.10}$ $74.40^{0.30}$ $74.90^{0.10}$ $66.70^{5.10}$ $74.70^{2.20}$ 14.57
GraphSAGE $79.70^{1.70}$ $72.40^{3.60}$ $49.90^{5.00}$ $65.80^{4.90}$ $63.90^{7.70}$ $65.90^{2.70}$ $64.67^{2.41}$ 14.67
DGCNN $73.76^{0.49}$ $70.03^{0.86}$ $47.83^{0.85}$ $79.37^{0.94}$ $74.44^{0.47}$ $58.59^{2.47}$ $75.54^{0.94}$ 14.00
CapsGNN $79.62^{0.91}$ $73.10^{4.83}$ $50.27^{2.65}$ $75.38^{4.17}$ $78.35^{1.55}$ $66.00^{1.80}$ $76.28^{3.63}$ 9.57
GIN $80.20^{1.90}$ $75.10^{5.10}$ $52.30^{2.80}$ $75.20^{2.90}$ $79.10^{1.40}$ $64.60^{7.00}$ $76.20^{2.80}$ $68.44^{1.89}$ 7.75
GDGIN - - - $77.80^{3.60}$ $60.30^{4.50}$ $73.70^{3.40}$ 15.33
GraphMAE $80.32^{0.46}$ $75.52^{0.66}$ $51.63^{0.52}$ $80.40^{0.30}$ $75.30^{0.39}$ 6.00
InfoGraph $70.65^{1.13}$ $73.03^{0.87}$ $49.69^{0.53}$ $72.85^{1.78}$ $73.80^{0.70}$ $61.65^{1.43}$ $74.40^{0.30}$ 15.14
JOAO $75.53^{0.18}$ $70.21^{3.08}$ $47.22^{0.41}$ $75.81^{0.73}$ $74.86^{0.39}$ $73.31^{0.48}$ 17.50
RGCL $70.92^{0.65}$ $71.85^{0.84}$ $78.86^{0.48}$ $75.03^{0.43}$ 15.75
CGKS $76.80^{0.10}$ $79.10^{0.20}$ $63.50^{1.30}$ $76.00^{0.20}$ 7.75
DIFFPOOL $75.48^{0.00}$ $73.14^{0.70}$ $49.53^{3.98}$ $80.64^{0.00}$ $62.32^{1.90}$ $76.25^{0.00}$ $61.98^{1.98}$ 11.00
SAGPool $78.85^{0.56}$ $72.55^{1.28}$ $49.33^{4.90}$ $76.45^{0.97}$ $74.18^{1.20}$ $71.86^{0.97}$ $74.06^{0.78}$ 14.71
ASAP $78.64^{0.50}$ $72.81^{0.50}$ $50.78^{0.75}$ $76.87^{0.70}$ $71.48^{0.42}$ $74.19^{0.79}$ $70.07^{0.55}$ 13.43
GMT $80.74^{0.54}$ $73.48^{0.76}$ $50.66^{0.82}$ $78.72^{0.59}$ $75.09^{0.59}$ 9.60
SAEPool - - - - $74.48^{0.00}$ $80.36^{0.00}$ $75.85^{0.00}$ 6.67
MVPool - $52.00^{0.80}$ $82.20^{1.40}$ $80.10^{1.30}$ $\underline{\bf{85.70}}^{1.20}$ $81.90^{1.60}$ 6.20
PAS - - $53.13^{4.49}$ $79.62^{1.75}$ $77.36^{3.69}$ 5.33
GraphGPS - - - - $79.44^{0.65}$ $53.75^{6.20}$ $76.27^{0.95}$ 12.67
U2GNN $77.84^{1.48}$ $77.04^{3.45}$ $53.60^{3.53}$ $80.23^{1.48}$ $69.63^{3.60}$ $78.53^{4.07}$ 4.17
UGT - - - - $77.55^{0.16}$ $80.12^{0.32}$ $75.45^{1.26}$ 6.33
CGIPool $80.30^{0.69}$ $72.40^{0.87}$ $51.45^{0.65}$ $78.62^{1.04}$ $74.10^{2.31}$ $77.94^{1.37}$ 9.83
DosCond - - - $78.92^{0.64}$ $71.70^{0.20}$ 13.00
KGC - $69.20^{1.37}$ $61.58^{2.49}$ $66.43^{0.92}$ 18.33
SS - $73.89^{0.70}$ $51.80^{1.20}$ $79.78^{0.49}$ $76.13^{0.60}$ 6.00
GRLsc $\underline{\bf{98.18}}^{0.42}$ $\underline{\bf{89.40}}^{2.50}$ $\underline{\bf{63.20}}^{3.55}$ $\underline{\bf{99.58}}^{0.78}$ $79.08^{1.92}$ $\underline{\bf{76.43}}^{5.22}$ $76.37^{4.37}$ $76.98^{1.90}$ 2.63
5.1.2. Baselines.
@@
To fully evaluate the effectiveness of our model, we select 28 related works from 6 categories for comparison.
@@
As for the Graph Transformer framework, we pick 17 baselines. They are (I) 2 kernel-based methods: GK (Shervashidze et al., 2009) and WL (Shervashidze et al., 2011); (II) 8 GNN-based methods: GCN (Kipf and Welling, 2017), GAT (Velickovic et al., 2018), GraphSAGE (Hamilton et al., 2017), DGCNN (Zhang et al., 2018), CapsGNN (Xinyi and Chen, 2019), GIN (Xu et al., 2019), GDGIN (Kong et al., 2022), and GraphMAE (Hou et al., 2022); (III) 4 Contrastive Learning methods: InfoGraph (Sun et al., 2020), JOAO (You et al., 2021), RGCL (Shuai et al., 2022), and CGKS (Zhang et al., 2023b); and (IV) 3 GT-based methods: GraphGPS (Rampásek et al., 2022), U2GNN (Nguyen et al., 2022), and UGT(Hoang and Lee, 2024).
@@
We also choose 11 models from 2 categories to verify our Graph Coarsening technique. One is (V) 7 Graph Pooling methods: DIFFPOOL (Ying et al., 2018), SAGPool(Lee et al., 2019), ASAP (Ranjan et al., 2020), GMT (Baek et al., 2021), SAEPool (Zhu et al., 2022), MVPool (Zhang et al., 2023a), and PAS (Wei et al., 2024). The other is (VI) 4 Graph Coarsening and Condensing methods: CGIPool (Pang et al., 2021), DosCond (Jin et al., 2022a), KGC (Chen et al., 2023c), and SS (Zhang and Zhao, 2024).
@@
To further explain the unique design of our LCC block, we select 3 coarsening techniques: networkx (Hagberg et al., 2008), KGC (Chen et al., 2023c), and L19 (Loukas, 2019). Each one takes two variants: neighborhood and clique. We will explain the details later in Section 5.4.
5.1.3. Implementation Details.
@@
We follow the work (Xu et al., 2019; Nguyen et al., 2022) to evaluate the performance of our proposed method, adopting accuracy as the evaluation metric for graph classification tasks. To ensure a fair comparison between methods, we use the same data splits and 10-fold cross-validation accuracy to report the performance. In detail, we set the batch size to 4 and the dropout to 0.5. We pick the hierarchy depth in $\{1,2,3\}$, the number of Transformer Encoder layers in $\{1,2,3,4\}$, the number of sampling neighbors in $\{4,8,16\}$, the hidden size in $\{128,256,512,1024\}$, and the initial learning rate in $\{5e^{-4},1e^{-4},5e^{-4},1e^{-3}\}$. We utilize Adam (Kingma and Ba, 2015) as the optimizer. All experiments are trained and evaluated on an NVIDIA RTX 3050 OEM 8GB GPU and 16GB CPU. Our code is available: https://github.com/NickSkyyy/LCC-for-GC.
5.2. Results and Analysis
5.2.1. Main Results.
@@ 5.2. Results and Analysis
Table 2 shows the main results of the graph classification task. GRLsc outperforms other baseline models on most datasets, with the highest average ranking 2.63 above all. In detail, the boost of GRLsc is higher in social networks than in molecules and bioinformatics. WL, GraphMAE, CGKS, PAS, U2GNN, and SS show competitive results among their categories, even better than GRLsc (e.g., WL in N1 and N109). The main reason is that cliques of social networks are more distinct, while plain cliques and loop structures can not cover functional molecules and proteins completely. These main experimental results show the limitation of our model: the classification ability slightly drops when fewer loops and cliques appear. N1 and N109 contain many simple loops and long chains, while some unconnected graphs in PRO make it hard to coarsen and convert. We will show more details later in the ablation of the LCC block (Section 5.4) and case study (Section 5.6).
5.2.2. Robustness Studies.
We also investigate the robustness of GRLsc. Figure  5 presents the model performance of 10-fold in the best epoch for each dataset. Most experiment folds are higher than or equal to the average, having a mean turbulence of 2.58%. The worst two are PTC and PRO, with standard deviations of 5.22% and 4.37%, respectively. In PTC, the performance gap between the best fold (1 and 4, 82.86%) and the worst fold (5 and 9, 67.65%) is 15.21%, about 5 in 34 graphs. The best two are CO and DD, with standard deviations of 0.42% and 0.78%, respectively. In CO, the performance gap between the best fold (1, 98.80%) and the worst fold (8, 97.20%) is 1.60%, about 8 in 500 graphs.
5.3. Block Ablation Study
In this section, we will evaluate the effectiveness of each component of our model. We design two sets of ablation experiments, one for modules to analyze the effectiveness of GRLsc. And the other to discuss the impact of different coarsening strategies on downstream classification tasks.
@@ 5.3. Block Ablation Study
Figure 5. 10 folds / best epoch on every dataset. The red dashed line plots the average accuracy.
Table 3. Block ablation studies on all eight datasets.
Method CO IB IM DD N1 PTC PRO N109
GRLsc $\underline{\bf{98.18}}^{0.42}$ $\underline{\bf{89.40}}^{2.50}$ $\underline{\bf{63.20}}^{3.55}$ $\underline{\bf{99.58}}^{0.78}$ $\underline{\bf{79.08}}^{1.92}$ $\underline{\bf{76.43}}^{5.22}$ $\underline{\bf{76.37}}^{4.37}$ $\underline{\bf{76.98}}^{1.90}$
- ${}_{w/o\,LGC}$ $97.96^{0.56}$ $87.60^{2.58}$ $61.47^{3.10}$ $99.32^{1.06}$ $77.42^{1.64}$ $69.50^{6.24}$ $75.02^{4.12}$ $74.90^{2.05}$
- ${}_{w/o\,LCC}$ $97.60^{0.81}$ $87.10^{2.91}$ $58.60^{3.74}$ $83.78^{4.63}$ $75.33^{2.12}$ $66.92^{8.38}$ $74.48^{3.42}$ $72.79^{2.14}$
- ${}_{w/o\,LGC+LCC}$ $96.50^{0.84}$ $84.10^{2.91}$ $57.87^{3.98}$ $83.02^{5.33}$ $72.53^{2.28}$ $63.09^{7.90}$ $68.01^{3.13}$ $69.20^{1.04}$
5.4. Coarsening Strategy Ablation
To better discuss the impact of different coarsening strategies on the model performance, we select four categories of graph coarsening schemes: random, networkx (Hagberg et al., 2008), KGC (Chen et al., 2023c), and L19 (Loukas, 2019). Under each category, there are two variants: neighbor and clique.
  • neighbor is an algorithm based on adjacent structures, which mainly aggregates information according to 1-hop adjacent nodes. There is no such algorithm in networkx, so we pick nx.cycle instead. It is similar to clique but based on cycle searches.
  • clique is an algorithm based on various clique searches, which aggregate node information according to clique partitions.
Table 4. Coarsening strategy ablation studies on all eight datasets.
Method CO IB IM DD N1 PTC PRO N109
Neighbor $98.12^{0.65}$ $89.10^{4.09}$ $62.80^{3.45}$ $98.64^{1.94}$ $77.79^{1.48}$ $70.67^{6.53}$ $75.20^{3.99}$ $76.40^{1.99}$
Random $97.94^{0.58}$ $89.10^{3.78}$ $61.70^{1.90}$ $99.32^{0.74}$ $77.79^{2.36}$ $71.81^{4.61}$ $72.51^{5.68}$ $74.10^{1.80}$
nx.cycle $98.02^{0.60}$ $87.00^{2.31}$ $62.47^{4.06}$ $86.16^{4.10}$ $74.84^{2.59}$ $75.92^{6.31}$ $74.30^{4.57}$ $73.81^{2.21}$
nx.clique $98.04^{0.58}$ $87.90^{2.62}$ $62.53^{3.41}$ $86.19^{5.62}$ $75.69^{1.92}$ $72.05^{6.67}$ $72.77^{6.16}$ $73.88^{1.81}$
KGC.neighbor $98.08^{0.58}$ $87.00^{1.95}$ $58.87^{4.03}$ $97.71^{2.57}$ $76.18^{2.12}$ $75.00^{2.66}$ $73.13^{3.70}$ $73.42^{1.36}$
KGC.clique $96.06^{0.77}$ $89.30^{2.72}$ $59.40^{3.66}$ $82.00^{7.04}$ $74.60^{1.91}$ $70.61^{8.86}$ $74.75^{5.05}$ $72.84^{1.72}$
L19.neighbor $96.92^{0.84}$ $87.10^{2.62}$ $61.70^{3.17}$ $96.44^{4.82}$ $75.94^{2.20}$ $74.06^{8.07}$ $74.21^{4.45}$ $73.81^{2.14}$
L19.clique $97.28^{0.66}$ $88.20^{2.79}$ $61.87^{3.55}$ $91.08^{3.00}$ $77.81^{1.48}$ $71.20^{6.10}$ $74.30^{3.65}$ $75.14^{1.86}$
LCC $\underline{\bf{98.18}}^{0.42}$ $\underline{\bf{89.40}}^{2.50}$ $\underline{\bf{63.20}}^{3.55}$ $\underline{\bf{99.58}}^{0.78}$ $\underline{\textbf{79.08}}^{1.92}$ $\underline{\bf{76.43}}^{5.22}$ $\underline{\textbf{76.37}}^{4.37}$ $\underline{\textbf{76.98}}^{1.90}$
In terms of implementation, we set $K=5$ and $r=0.5$, according to the paper (Chen et al., 2023c; Loukas, 2019). The algorithm strategy is greedy, and other parameters are consistent with the origin. To ensure alignment with GRLsc, the maximum coarsening level is set to 1, which means only coarsening the original graph without higher-level compressions. The random algorithm sets $R=K=5$, and each supernode randomly selects components from the sampling sets until all nodes are covered. Table 4 shows the model performance under different coarsening strategies, where LCC achieves the optimal results under all eight datasets. It indicates that the graph coarsening view obtained by LCC is more suitable for graph classification tasks.
5.4.1. Time Analysis

### v5

@@ 5.1.1 Datasets
We evaluate our approach on eight datasets from TUDataset [47], including three social network datasets: COLLAB (CO), IMDB-BINARY (IB), and IMDB-MULTI (IM), two molecules datasets: NCI1 (N1) and NCI109 (N109), and three bioinformatics datasets: D&D (DD), PTC_MR (PTC), and PROTEINS (PRO). Table 1 shows the details of the datasets.
5.1.2 Baselines
@@
To fully evaluate the effectiveness of our model, we select 31 related works from 6 categories for comparison.
@@
As for the Graph Transformer framework, we pick 20 baselines. They are (I) 2 kernel-based methods: GK [48] and WL [49]; (II) 8 GNN-based methods: GCN [50], GAT [51], GraphSAGE [52], DGCNN [53], CapsGNN [54], GIN [41], GDGIN [55], and GraphMAE [56]; (III) 4 Contrastive Learning methods: InfoGraph [57], JOAO [58], RGCL [59], and CGKS [16]; and (IV) 6 GT-based methods: GKAT [60], GraphGPS [1], U2GNN [37], SAN-SSCA [61], SA-GAT [2], and UGT[3].
@@
We also choose 11 models from 2 categories to verify our Graph Coarsening technique. One is (V) 7 Graph Pooling methods: DIFFPOOL [9], SAGPool[25], ASAP [62], GMT [14], SAEPool [63], MVPool [26], and PAS [64]. The other is (VI) 4 Graph Coarsening and Condensing methods: CGIPool [27], DosCond [22], KGC [28], and SS [29].
@@
To further explain the unique design of our LCC, we select 3 coarsening techniques: networkx [65], KGC [28], and L19 [30]. Each one takes two variants: neighborhood and clique. We will explain the details later in Section 5.5.
5.1.3 Implementation Details
@@
We follow the work [41, 37] to evaluate the performance of our proposed method, adopting accuracy as the evaluation metric for graph classification tasks. To ensure a fair comparison between methods, we use the same data splits and 10-fold cross-validation accuracy to report the performance. All experiments are trained and evaluated on an NVIDIA RTX 3050 OEM 8GB GPU. Our code is available: https://github.com/NickSkyyy/LCC-for-GC.
5.2 Graph Classification Results
Table 2: The main results of the graph classification task on four of eight datasets (mean accuracy (%) and standard deviation). The best scores are in bold and underline. ’-’ indicates that results are unavailable in the original or published papers. A.R shows the average ranking of each method. We highlight every method with outstanding performance in each category.
Method CO DD PTC PRO A.R
GW $72.84^{0.28}$ $78.45^{0.26}$ $57.26^{1.41}$ $71.67^{0.55}$ 21.00
WL $79.02^{1.77}$ $79.78^{0.36}$ $57.97^{0.49}$ $74.68^{0.49}$ 9.63
GCN $81.72^{1.64}$ $79.12^{3.07}$ $59.40^{10.30}$ $75.65^{3.24}$ 10.25
GAT $75.80^{1.60}$ $74.40^{0.30}$ $66.70^{5.10}$ $74.70^{2.20}$ 16.14
GraphSAGE $79.70^{1.70}$ $65.80^{4.90}$ $63.90^{7.70}$ $65.90^{2.70}$ 15.13
DGCNN $73.76^{0.49}$ $79.37^{0.94}$ $58.59^{2.47}$ $75.54^{0.94}$ 16.14
CapsGNN $79.62^{0.91}$ $75.38^{4.17}$ $66.00^{1.80}$ $76.28^{3.63}$ 10.71
GIN $80.20^{1.90}$ $75.20^{2.90}$ $64.60^{7.00}$ $76.20^{2.80}$ 8.88
GDGIN - $77.80^{3.60}$ $60.30^{4.50}$ $73.70^{3.40}$ 17.33
GraphMAE $80.32^{0.46}$ $75.30^{0.39}$ 6.60
InfoGraph $70.65^{1.13}$ $72.85^{1.78}$ $61.65^{1.43}$ $74.40^{0.30}$ 17.43
JOAO $75.53^{0.18}$ $75.81^{0.73}$ $73.31^{0.48}$ 19.50
RGCL $70.92^{0.65}$ $78.86^{0.48}$ $75.03^{0.43}$ 17.25
CGKS $76.80^{0.10}$ $63.50^{1.30}$ $76.00^{0.20}$ 10.00
DIFFPOOL $75.48^{0.00}$ $80.64^{0.00}$ $76.25^{0.00}$ 13.14
SAGPool $78.85^{0.56}$ $76.45^{0.97}$ $71.86^{0.97}$ 16.14
ASAP $78.64^{0.50}$ $76.87^{0.70}$ $74.19^{0.79}$ 15.14
GMT $80.74^{0.54}$ $78.72^{0.59}$ $75.09^{0.59}$ 10.20
SAEPool - - - $80.36^{0.00}$ 9.00
MVPool - $82.20^{1.40}$ $\underline{\bf{85.70}}^{1.20}$ 7.20
PAS - $79.62^{1.75}$ $77.36^{3.69}$ 6.00
GKAT $73.10^{2.00}$ $78.60^{3.40}$ $75.80^{3.80}$ 16.50
GraphGPS - - - $53.75^{6.20}$ 13.67
U2GNN $77.84^{1.48}$ $80.23^{1.48}$ $69.63^{3.60}$ $78.53^{4.07}$ 4.67
SAN-SSCA $73.96^{0.00}$ $82.41^{0.00}$ 5.50
SA-GAT - $91.40^{0.55}$ $66.93^{1.49}$ $77.80^{0.58}$ 3.80
UGT - - - $80.12^{0.32}$ 7.67
CGIPool $80.30^{0.69}$ $74.10^{2.31}$ 10.67
DosCond - $78.92^{0.64}$ 17.00
KGC - - $61.58^{2.49}$ $66.43^{0.92}$ 20.00
SS - $79.78^{0.49}$ $76.13^{0.60}$ 7.25
LCC4GC $\underline{\bf{98.18}}^{0.42}$ $\underline{\bf{99.58}}^{0.78}$ $\underline{\bf{76.43}}^{5.22}$ $76.37^{4.37}$ 3.25
Table 3: The rest four datasets due to limited space.
Method IB IM N1 N109 A.R
GW $65.87^{0.98}$ $43.89^{0.38}$ $62.30^{0.30}$ 21.00
WL $73.40^{4.63}$ $49.33^{4.75}$ $82.19^{0.18}$ $82.46^{0.24}$ 9.63
GCN $73.30^{5.29}$ $51.20^{5.13}$ $76.00^{0.90}$ $67.09^{3.43}$ 10.25
GAT $70.50^{2.30}$ $47.80^{3.10}$ $74.90^{0.10}$ 16.14
GraphSAGE $72.40^{3.60}$ $49.90^{5.00}$ $64.67^{2.41}$ 15.13
DGCNN $70.03^{0.86}$ $47.83^{0.85}$ $74.44^{0.47}$ 16.14
CapsGNN $73.10^{4.83}$ $50.27^{2.65}$ $78.35^{1.55}$ 10.71
GIN $75.10^{5.10}$ $52.30^{2.80}$ $79.10^{1.40}$ $68.44^{1.89}$ 8.88
GraphMAE $75.52^{0.66}$ $51.63^{0.52}$ $80.40^{0.30}$ 6.60
InfoGraph $73.03^{0.87}$ $49.69^{0.53}$ $73.80^{0.70}$ 17.43
JOAO $70.21^{3.08}$ $47.22^{0.41}$ $74.86^{0.39}$ 19.50
RGCL $71.85^{0.84}$ 17.25
CGKS - - $79.10^{0.20}$ 10.00
DIFFPOOL $73.14^{0.70}$ $49.53^{3.98}$ $62.32^{1.90}$ $61.98^{1.98}$ 13.14
SAGPool $72.55^{1.28}$ $49.33^{4.90}$ $74.18^{1.20}$ $74.06^{0.78}$ 16.14
ASAP $72.81^{0.50}$ $50.78^{0.75}$ $71.48^{0.42}$ $70.07^{0.55}$ 15.14
GMT $73.48^{0.76}$ $50.66^{0.82}$ 10.20
SAEPool - - $74.48^{0.00}$ $75.85^{0.00}$ 9.00
MVPool $52.00^{0.80}$ $80.10^{1.30}$ $81.90^{1.60}$ 7.20
PAS - $53.13^{4.49}$ 6.00
GKAT $71.40^{2.60}$ $47.50^{4.50}$ $75.20^{2.40}$ 16.50
GraphGPS - - $79.44^{0.65}$ $76.27^{0.95}$ 13.67
U2GNN $77.04^{3.45}$ $53.60^{3.53}$ 4.67
SAN-SSCA - - $\underline{\textbf{84.92}}^{0.00}$ $\underline{\textbf{82.82}}^{0.00}$ 5.50
SA-GAT $75.59^{0.77}$ $79.32^{1.09}$ 3.80
UGT - - $77.55^{0.16}$ $75.45^{1.26}$ 7.67
CGIPool $72.40^{0.87}$ $51.45^{0.65}$ $78.62^{1.04}$ $77.94^{1.37}$ 10.67
DosCond - - $71.70^{0.20}$ 17.00
KGC $69.20^{1.37}$ 20.00
SS $73.89^{0.70}$ $51.80^{1.20}$ 7.25
LCC4GC $\underline{\bf{89.40}}^{2.50}$ $\underline{\bf{63.20}}^{3.55}$ $79.08^{1.92}$ $76.98^{1.90}$ 3.25
@@ 5.2 Graph Classification Results
Tables 2 and  3 show the main results of the graph classification task. LCC4GC outperforms other baseline models on most datasets, with the highest average ranking 3.25 above all. In detail, the boost of LCC4GC is higher in social networks than in molecules and bioinformatics. WL, GraphMAE, CGKS, PAS, SA-GAT, and SS show competitive results among their categories, even better than LCC4GC (e.g., WL in N1 and N109). The main reason is that cliques of social networks are more distinct, while plain cliques and loop structures can not cover functional molecules and proteins completely. These main experimental results show the limitation of our model: the classification ability slightly drops when fewer loops and cliques appear. N1 and N109 contain many simple loops and long chains, while some unconnected graphs in PRO make it hard to coarsen and convert. We will show more details later in the ablation of the LCC (Section 5.5) and case study (Section 5.6).
5.3 Block Ablation Study
Table 4: Block ablation studies on all eight datasets.
Method CO DD PTC PRO
LCC4GC $\underline{\bf{98.18}}^{0.42}$ $\underline{\bf{99.58}}^{0.78}$ $\underline{\bf{76.43}}^{5.22}$ $\underline{\bf{76.37}}^{4.37}$
${}_{w/o\,LGC}$ $97.96^{0.56}$ $99.32^{1.06}$ $69.50^{6.24}$ $75.02^{4.12}$
${}_{w/o\,LCC}$ $97.60^{0.81}$ $83.78^{4.63}$ $66.92^{8.38}$ $74.48^{3.42}$
${}_{w/o\,both}$ $96.50^{0.84}$ $83.02^{5.33}$ $63.09^{7.90}$ $68.01^{3.13}$
Method IB IM N1 N109
LCC4GC $\underline{\bf{89.40}}^{2.50}$ $\underline{\bf{63.20}}^{3.55}$ $\underline{\bf{79.08}}^{1.92}$ $\underline{\bf{76.98}}^{1.90}$
${}_{w/o\,LGC}$ $87.60^{2.58}$ $61.47^{3.10}$ $77.42^{1.64}$ $74.90^{2.05}$
${}_{w/o\,LCC}$ $87.10^{2.91}$ $58.60^{3.74}$ $75.33^{2.12}$ $72.79^{2.14}$
${}_{w/o\,both}$ $84.10^{2.91}$ $57.87^{3.98}$ $72.53^{2.28}$ $69.20^{1.04}$
@@ 5.3 Block Ablation Study
5.4 Hyper-parameters Study
As for $K$ and $T$, a deeper and more complicated model improves model performances and enhances the ability to capture complex structural information. However, as the hierarchy extends continuously, overfitting leads to a decline in classification accuracy. The most striking example is the plot of CO and DD on $T$, where the performance of LCC4GC decreases substantially when the number of layers changes from 3 to 4.
Table 5: Optimal parameter settings. We set $|H|=1024$.
Parameters CO DD PTC PRO IB IM N1 N109
$K$ 1 2 3 2 1 3 1 1
$T$ 2 2 4 1 2 1 2 1
$N$ 16 16 16 4 16 4 16 8
Figure 4: Hyper-parameters Study. We evaluate four hyper-parameters: hierarchy depth $K$, the number of Transformer Encoder layers $T$, the number of sampling neighbors $N$, and the hidden size $|H|$.